ABCD is square and forces acting at a point are represented in magnitude and direction by AB. 2BC. 2CD And DA. Find the resultant

Grade:12th pass

1 Answers

Harsh Tiwari
2 years ago
On square forces will be equal in magnitude.
 
Resultant = AB +2 BC + 2CD + DA 
Let AB = p 
Then
BC = p vector
 
CD = -p vector
Da = -p vector
 
Then
Resultant = p + 2p -2p -p=0
 
Hence resultant will be 0